OpenStudy (rebeccapink6398):

@31356 could you help me? :D What is the value of x?

OpenStudy (31356):

Problem with it?

OpenStudy (rebeccapink6398):

x/8=9/24

OpenStudy (31356):

Thx for the medal :D

OpenStudy (rebeccapink6398):

:D yup!... heres the options :D A. 3 B. 6 C. 27 D. 72

OpenStudy (31356):

It's the first one :D

OpenStudy (rebeccapink6398):

A?

OpenStudy (31356):

Since 9/24= 0.375 3/8 is 0.375 too :D

OpenStudy (rebeccapink6398):

:D thx!

OpenStudy (31356):

My pleasure :D

OpenStudy (31356):

Yea, just plug in the variables :D

OpenStudy (rebeccapink6398):

awe :( I got 40%..u got the problem right tho lol

OpenStudy (rebeccapink6398):

could you help me re-take it?

OpenStudy (rebeccapink6398):

my mo was helping me and we both don't get it

OpenStudy (rebeccapink6398):

mom*

OpenStudy (rebeccapink6398):

@31356 ?

OpenStudy (31356):

Oh sorry, I was doing my chores

OpenStudy (rebeccapink6398):

lol sorry :)

OpenStudy (31356):

At least you didn't made a 0% though, lol :D

OpenStudy (31356):

That's okay :D

OpenStudy (rebeccapink6398):

LOL! :D

OpenStudy (31356):

Yea, Triple smileys :D:D:D

OpenStudy (rebeccapink6398):

lol :D :D :D could you help me re-take the test>?

OpenStudy (31356):

Of course I can help you :D

OpenStudy (rebeccapink6398):

:D

OpenStudy (31356):

Questions?

OpenStudy (rebeccapink6398):

1.State whether the ratios form a proportion 1/4 and 2/5

OpenStudy (31356):

What do you mean by that? Lol ,confused too

OpenStudy (rebeccapink6398):

2.Solve for the variable 3/9 = p/24

OpenStudy (rebeccapink6398):

not sure that's what they said... oh and I was saying 1. for first question and 2. for second question

OpenStudy (31356):

Oh okay

OpenStudy (rebeccapink6398):

..r u still confused?

OpenStudy (31356):

I am going to ask someone, #2 is easy

OpenStudy (rebeccapink6398):

lol ok brb

OpenStudy (31356):

#2 I got 8

OpenStudy (31356):

24*3=72 72/9=8

OpenStudy (31356):

Is #1 a yes or no question?

OpenStudy (rebeccapink6398):

Yes it is :)

OpenStudy (31356):

Oh wow, perfect guess for me, lol :D

OpenStudy (31356):

Do you remember what you put it as at first?

OpenStudy (rebeccapink6398):

umm... they change the questions.

OpenStudy (31356):

Oh wait I get it, it is NO for the 1st problem! :D

OpenStudy (rebeccapink6398):

ok :D

OpenStudy (31356):

Cross multiply 5 and 8 is not the same! :D

OpenStudy (rebeccapink6398):

here is the questions :) 1.State whether the ratios form a proportion 1/4 and 2/5~DONE 2.Solve for the variable 3/9 = p/24~DONE 3.Solve for the variable k/3 = 16/24 4.The ratio of red marbles to blue marbles in a jar is 3 to 8. There are 40 blue marbles in the jar. How many red marbles are in the jar? 5.What is the value of x? x/12=55/60

OpenStudy (31356):

#3 = 2

OpenStudy (rebeccapink6398):

ok they have that.. :D

OpenStudy (31356):

#4 = 15

OpenStudy (31356):

#5 = 11

OpenStudy (31356):

That was easy, lol :D

OpenStudy (31356):

Score? o.0

OpenStudy (rebeccapink6398):

100%!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!! :D :D :D :D :D :D :D :D :D :D SEVERAL SMILYS!!!! lol!

OpenStudy (rebeccapink6398):

:D?

OpenStudy (31356):

Cool! :D

OpenStudy (31356):

Medal for you! :D

OpenStudy (rebeccapink6398):

:D lol..u give me medal? why? lol.. u did all the work :( which makes me feel bad...

OpenStudy (rebeccapink6398):

what grade r u in?

OpenStudy (31356):

You deserved it, lol :D

OpenStudy (31356):

8th grade and you?

OpenStudy (rebeccapink6398):

6..no wonder its easy 4 u! lol

OpenStudy (31356):

Yea lol :D

OpenStudy (31356):

Yea, I am learning Algebra now so, yea it's easy! :D